OpenStudy (anonymous):

Simplify (10 + 3i)(5 – 2i)

OpenStudy (angela210793):

50-20i+15i+6=56-5i

OpenStudy (saifoo.khan):

Agree^

OpenStudy (lalaly):

50 -20i +15i +6 = 56-5i

OpenStudy (saifoo.khan):

Agree^

OpenStudy (anonymous):

Agree^

OpenStudy (saifoo.khan):

please dont copy me @solver. jk